OpenStudy (trisarahtops):

Help Please! Find the values of m and b that make the following function differentiable. (This is a piecewise function.) f(x)= {x^3, x≤1 {mx+b, x>1

jimthompson5910 (jim_thompson5910):

what is the value of x^3 when x = 1?

OpenStudy (trisarahtops):

3

jimthompson5910 (jim_thompson5910):

x^3 doesn't mean x times 3

OpenStudy (trisarahtops):

1 lol

OpenStudy (trisarahtops):

srry I didn't read that right

jimthompson5910 (jim_thompson5910):

yep x^3 is equal to 1 when x = 1

jimthompson5910 (jim_thompson5910):

what is mx+b equal to when x = 1 ?

OpenStudy (trisarahtops):

1=m(1)+b

jimthompson5910 (jim_thompson5910):

so you should find that 1 = m+b solve for m or b. I'm going to solve for b to get b = 1-m

jimthompson5910 (jim_thompson5910):

now let's derive each piece x^3 derives to 3x^2 mx+b derives to just m agreed? or no?

OpenStudy (trisarahtops):

Yes agreed

jimthompson5910 (jim_thompson5910):

ok for the function f(x) to be differentiable at x = 1, f ' (x) needs to be continuous at x = 1 so we need to plug x = 1 into each piece (3x^2 and m) and set those outputs equal to each other first piece: 3x^2 ----> 3(1)^2 = 3 second piece: m ----> m ... there are no x values to plug in x = 1 so we know that m = 3

OpenStudy (trisarahtops):

so it's 3x^2= 3(1)+b so far

jimthompson5910 (jim_thompson5910):

if m = 3 and we know b = 1-m, then what is b?

OpenStudy (trisarahtops):

-2=b

jimthompson5910 (jim_thompson5910):

good

OpenStudy (trisarahtops):

Thank you :)

jimthompson5910 (jim_thompson5910):

no problem